OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 – 6x – 7. Find f(g(2)).

OpenStudy (anonymous):

g(2)=2^2-6x2-7=-15

OpenStudy (anonymous):

f(g(2))=f(-15)=-15+8=-7

OpenStudy (anonymous):

wait so what do i do from there?